A 20.0-kg horizontal plank 4.00 m long rests on two supports, one at the left end and a second 1 .00 m right end. What is the magnitude of the force exerted on the plank by the support near the right end?
1
Expert's answer
2020-12-06T17:32:26-0500

"L=4m
\\mass \space of plank\space, m=20kg
\\ take \space counterclockwise \space torques\space are positive\\
Apply\space torque \space equation\space about\space one\space end\\
F_{1}(0)-mg(\frac{L}{2})+F_{2}(L-1m)=0
\\F_{2}(L-1m)=mg(\frac{L}{2})
\\F_{2}=mg(\frac{L}{2(L-1)})\\
\quad=20kg\times 9.8m/s^{2}\times\frac{4m}{2(4m-1m)})\\
\quad=130.66N\\
\quad=131N"